# Deep Active Speech Cancellation with Multi-Band Mamba Network

### v1

@@ 3.2 Optimization Objective
4 Experiments and Results
Datasets: The training data is sourced from the AudioSet dataset (Gemmeke et al., 2017), which we encompassed 248 distinct audio categories. These categories include various types of ambient sounds such as hubbub, speech noise, and speech babble. The dataset comprises 22,224 audio samples, totaling 18.5 hours of audio content. To maintain consistency with ARN method (Zhang et al., 2023a), each audio sample was standardized to a duration of 3 seconds and resampled to a 16kHz. Additionally, 20,000 samples (90%) of the dataset were allocated for training. The remaining 2,224 samples were reserved for testing. The test sets were obtained from the NOISEX dataset (Varga & Steeneken, 1993), which includes a wide range of noise types, such as bubble noise, factory noise, and engine noise. Additionally, we utilized the test sets from the following speech datasets: TIMIT (Garofolo, 1993), which contains recordings from 24 speakers representing 8 dialect regions; LibriSpeech (Panayotov et al., 2015), which includes 40 speakers from audiobook readings; and the Wall Street Journal (WSJ) (Garofolo et al., 1993), which features 8 speakers reading news articles.
[…]
Hyperparameters: An extensive grid search and cross-validation were employed to determine the optimal hyperparameters for each method. The hyperparameter values reported here correspond to the configurations that achieved the best performance in our experimental setup. The DeepASC architecture was trained using multiple numbers of subbands $Q$, specifically $Q=0$ (a single full band), $2$ and $3$. The bands decomposition filters are generated using the $scipy.signal.firwin$ function and applied to the signal via $torch.conv1d$. The temporal duration $M$ was set to 48,000 samples, corresponding to 3-second audio signals sampled at 16 kHz. The channel dimension $C$ was set to 256, and the kernel size $W$ was defined as 16. A batch size of 2 was used for training the DeepASC architecture. The Adam optimizer (Diederik, 2014) was employed with an initial learning rate of $1.5\times 10^{-4}$. A learning rate decay factor of 0.5 was applied every 2 epochs after an initial warm-up period of 30 epochs. Gradient clipping with a threshold of 5 was applied to prevent exploding gradients.
Baseline Methods: We compared our proposed method with several established ANC techniques, including Deep ANC (Zhang & Wang, 2021), Attentive Recurrent Network (ARN) (Zhang et al., 2023a), Filtered-x Least Mean Squares (FxLMS), and Tangent Hyperbolic Function FxLMS (THF-FxLMS), (Ghasemi et al., 2016). All methods were evaluated in both linear and nonlinear simulations, considering both noise and speech signals. FxLMS, Deep ANC, and ARN were implemented and trained by us. All methods were evaluated under identical simulation conditions. For the learned methods, namely Deep ANC and ARN, the same training dataset used for our proposed method was applied, and we ensured the reproduction of results consistent with those reported in the respective papers. In our Deep ANC implementation, we employed 20-ms short-time Fourier transform (STFT) frames with a 10-ms overlap between consecutive frames. For ARN, we utilized 16-ms frames with an 8-ms overlap. These baseline methods were selected to provide a comprehensive comparison across various ANC paradigms, encompassing both traditional adaptive filtering techniques and more recent deep learning approaches.
4.1 Noise Cancellation
Table 2 presents the NMSE for ANC algorithms across three noise types—engine, factory, and babble—using 3-second signal segments extracted from the NOISEX-92 dataset. For each noise type, the models were evaluated both without nonlinear distortions (where $\eta^{2}=\infty$) and with nonlinear distortions at $\eta^{2}=0.1$ and $\eta^{2}=0.5$. In the case of non-deep learning-based methods, namely FxLMS and THF-FxLMS, gradient clipping at $41e-$ 4 was applied due to the sensitivity of these algorithms to the step size, which caused instability during validation. The step sizes for these methods were set to 0.05 for engine noise, 0.4 for factory noise, and 0.3 for babble noise. The results indicate that these algorithms perform suboptimally compared to deep learning-based approaches.
Among the deep learning-based methods, and without considering the nonlinearity saturation effect, the proposed DeepASC method achieves state-of-the-art results. Specifically, for the case where $\eta^{2}=\infty$ it improves performance over the ARN method by 4.29 dB, 4.64 dB, and 7.26 dB for engine, factory, and babble noise, respectively. In the presence of nonlinear distortions ($\eta^{2}=0.5$), DeepASC continues to outperform ARN, with improvements of 4.36 dB, 4.62 dB, and 7.13 dB for engine, factory, and babble noise, respectively. For more severe nonlinearity ($\eta^{2}=0.1$), DeepASC still surpasses ARN with gains of 3.79 dB, 4.4 dB, and 5.76 dB. Figures 4(a), 4(b), and 4(c) offer visual comparisons of the different methods by plotting NMSE over time. These figures illustrate that the proposed DeepASC method consistently achieves superior NMSE performance compared to ARN, DeepANC, and FxLMS across almost every timestep.
@@ 4.1 Noise Cancellation
Table 3: Average NMSE for DeepASC and other algorithms across various speech datasets and nonlinear distortions.
@@
Table 4: Average NMSE, STOI & PESQ for ASC models in noisy speech with nonlinearity ($\eta^{2}=0.5$) and factory noise at different SNR.
Method Noise only SNR = 5dB SNR = 15dB SNR = 20dB
DeepANC -10.69 0.83 1.39 0.93 2.10 0.96 2.45
ARN -11.61 0.84 1.51 0.94 2.43 0.96 2.92
DeepASC -15.94 0.92 2.21 0.97 3.35 0.98 3.76
4.2 Speech Cancellation
Table 3 presents the average NMSE values for different ANC algorithms across three speech datasets: TIMIT, LibriSpeech, and WSJ, with speech segments affected by varying levels of nonlinear distortions. It is evident that speech cancellation is a more challenging task compared to noise cancellation, as reflected in the performance degradation of the different algorithms.
As observed in the noise cancellation case, in speech cancellation, the non-deep learning methods—FxLMS and THF-FxLMS—demonstrate suboptimal performance compared to deep learning-based approaches. Among the deep learning methods, DeepASC achieves the best overall results, surpassing the other algorithms significantly.
In the case without nonlinear distortions ($\eta^{2}=\infty$), DeepASC shows improvements over ARN by 6.13 dB, 4.78 dB, and 5.95 dB for the TIMIT, LibriSpeech, and WSJ datasets, respectively. In the presence of moderate nonlinear distortions ($\eta^{2}=0.5$), DeepASC continues to outperform ARN, with improvements of 6.18 dB for TIMIT, 4.34 dB for LibriSpeech, and 5.99 dB for WSJ. Under more severe nonlinear distortions ($\eta^{2}=0.1$), DeepASC maintains its superior performance, with enhancements of $5.97$dB, $2.46$dB, and $5.81$dB for TIMIT, LibriSpeech, and WSJ datasets, respectively. Figure 5 illustrates the performance of various ANC methods on a speech signal, comparing power spectra and spectrograms. DeepASC demonstrates superior noise suppression across all frequencies, including high frequencies, outperforming other methods such as DeepANC and ARN, which struggle more with high-frequency noise. Figure 4(d) further demonstrates DeepASC’s superior NMSE performance for speech signals, exceeding ARN, DeepANC, and FxLMS at nearly every time step.
4.3 Real-World Simulation
We expanded our investigation to assess the performance of our method in real-world settings, testing it across various simulation scenarios. This was necessary because the fixed task acoustic setup, which relies on the image method, has limitations regarding generalizability and real-world performance. We utilized the dataset from (Liebich et al., 2019), which includes acoustic paths from 23 individuals, measured in the real world and encompassing both primary and secondary paths. We applied DeepASC, along with baseline approaches, to the updated simulation conditions, evaluating their performance using Factory and Babble noise from the NoiseX-92 dataset and speech samples from the WSJ dataset. The results in Table 5 present the average NMSE across these categories. The results demonstrate that DeepASC consistently outperforms the alternative methods, achieving improvements of $2.80$dB in the Factory noise, $2.70$dB in the Babble noise, and $1.53$dB on the WSJ dataset.
Method/Dataset Factory ($\downarrow$) Babble ($\downarrow$) WSJ ($\downarrow$)
DeepANC -9.29 -10.94 -8.26
ARN -8.97 -11.17 -10.70
DeepASC -12.09 -13.87 -12.23
@@ 4.3 Real-World Simulation
Figure 5: Spectrograms and Power Spectra of Speech Signal (00da010c, WSJ) using Different ANC methods w/o nonlinearity ($\eta^{2}=\infty$)
Table 6: Average NMSE ($\downarrow$) of our method (w/o NOAS) for Noise and Speech using different number of bands, with $\eta^{2}=0.5$.
DeepASC (small) 1 -13.46 -14.26 -14.88 -13.22
DeepASC (medium) 1 -15.19 -15.82 -16.56 -14.86
DeepASC 3 -15.94 -16.36 -16.95 -15.32
DeepASC 4 -16.52 -16.55 -17.41 -15.84
4.4 Model Analysis
The number of frequency bands in the DeepASC architecture is a critical hyperparameter affecting performance. Table 6 compares DeepASC’s performance across different band configurations for the Factory noise, TIMIT, LibriSpeech, and WSJ datasets, with $\eta^{2}=0.5$. The ”1-band” models use a single full band, while the ”3-band” and ”4-band” models incorporate one medium band with two and three smaller sub-bands, respectively. A 2-band model, which would require two full bands, was excluded as it falls outside the intended design of DeepASC.
Table 7 compares model size and performance, with NMSE evaluated on factory noise under nonlinear distortion of $\eta=0.5$. DeepASC variants in this comparison are without NOAS optimization. The results indicate that even the smallest DeepASC configuration (1-band, small) outperforms the ARN architecture by $1.85$ dB, despite using only half the parameters ($8.0$M vs. $15.9$M). This is a significant outcome given the critical importance of model size in real-time ANC applications where latency is critical.
@@ 4.4 Model Analysis
The computational complexity of the models was assessed by comparing their FLOPs, averaged across 20 three-second samples from the Noisex-92 dataset, as presented in Table 8. The single-band, small variant of DeepASC demonstrated exceptional efficiency, requiring only 2.862G FLOPs while consistently surpassing the performance of the other models. This highlights its superior balance between computational cost and effectiveness. For further analysis, see Appendix B, which presents the Ablation Study.
Table 7: Comparison of ANC methods based on parameter size.
Models #Params NMSE ($\downarrow$)
DeepANC 8.8M -10.69
ARN 15.9M -11.61
DeepASC, 1 Band, S 8.0M -13.46
DeepASC, 1 Band, M 15.8M -15.19
DeepASC, 3 Bands 31.9M -15.94
DeepASC, 4 Bands 40.0M -16.52
Table 8: FLOPs & NMSE comparison for different ANC methods.
Method FLOPs (G) ($\downarrow$) NMSE ($\downarrow$)
DeepANC 7.199 -10.69
ARN 5.281 -11.61
DeepASC 2.419 -13.46
5 Conclusion
In this paper, we introduced a novel ASC approach using the Multi-Band Mamba architecture. By partitioning audio into frequency bands, our method enhances anti-signal generation and phase alignment. Combined with an optimization-driven loss function, it achieves near-optimal performance, improving both ANC and ASC outcomes. Our experimental results demonstrate a significant performance boost compared to state-of-the-art baselines, with improvements of 7.2dB in ANC and 6.2dB in ASC for voice audio signals. These results confirm the multi-band architecture’s effectiveness in handling diverse frequencies and real-world acoustic environments, where traditional methods often fail. Our approach addresses key challenges in the field by effectively leveraging frequency decomposition and optimization-based anti-signal generation, paving the way for more advanced audio cancellation technologies.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Overview of an ANC System
To elucidate the acoustic dynamics of ANC systems in in-ear headphones, Figure 6 presents a schematic representation of the Primary and Secondary acoustic paths.
The primary path $P(z)$ characterizes the transfer function between the external noise, as captured by the reference microphone, and the error microphone. This path models the propagation of ambient noise through the system. Conversely, the secondary path $S(z)$ represents the transfer function from the loudspeaker to the error microphone, encompassing the acoustic feedback loop within the ear canal. The schematic highlights the interaction between critical system components, including the processing unit, reference microphone, error microphone, and loudspeaker.
Appendix B Ablation Study
@@
To assess the contributions of the key components in the DeepASC architecture, we conducted an ablation study focusing on multiband processing, band size (small vs. medium), and the impact of NOAS optimization. Table 9 presents the results of this analysis, reporting the NMSE across four datasets: Factory, TIMIT, LibriSpeech, and WSJ, all evaluated under nonlinear distortion conditions ($\eta^{2}$ = 0.5).
Figure 6: Schematic representation of the acoustic paths in an in-ear ANC system.
@@
In our notation, ”+ S - Multiband - NOAS” refers to a small band configuration (8 mamba layers) without multiband processing or NOAS optimization, while ”+ S - Multiband + NOAS” refers to the same small band architecture with NOAS optimization applied. Similarly, ”+ M - Multiband - NOAS” represents a medium band configuration (16 mamba layers) without NOAS, and ”+ M - Multiband + NOAS” applies NOAS optimization to the same medium band model. The Full Method is defined as a configuration that employs one full medium band and two small sub-bands, with NOAS optimization applied.
@@
All models were initially trained using the ANC loss function defined in Eq. 10. Configurations with ”+ NOAS” were fine-tuned using NOAS optimization, whereas configurations with ”- NOAS” were trained exclusively using the ANC loss in Eq. 10. The results demonstrate that the removal of NOAS optimization consistently degrades performance across all datasets. For instance, on the Factory dataset, applying NOAS optimization to the small band model leads to a performance improvement of $0.73$dB, while the medium band model shows a larger improvement of $0.90$dB. This trend holds across the other datasets, reinforcing the crucial role of NOAS optimization in enhancing model performance. Multiband processing further improves the overall effectiveness of DeepASC. For example, the Full Method consistently outperforms the ”+ M - Multiband + NOAS” configuration, with gains of $0.14$dB, $0.2$dB, $0.16$dB, and $0.2$dB on the Factory, TIMIT, LibriSpeech, and WSJ datasets, respectively. Interestingly, the performance of the ”+ S - Multiband - NOAS” configuration is consistently lower than that of the ”+ M - Multiband - NOAS” variant across all datasets. Specifically, the small band model underperforms by $1.73$dB on Factory, $1.56$dB on TIMIT, $1.68$dB on LibriSpeech, and $1.66$dB on WSJ. This indicates that while multiband processing is valuable, the choice of band size plays a significant role in the model’s performance, with larger band sizes, particularly when combined with NOAS, yielding the best results.

### v2

@@ 6.1 Noise Cancellation
Table 6: Average NMSE (dB), STOI and PESQ for deep ANC models in noisy speech situations with LS nonlinearity ($\eta=0.5$) and factory noise at different SNR levels.
Method Noise only SNR = 5dB SNR = 15dB SNR = 20dB
DeepANC -10.69 0.83 1.39 0.93 2.10 0.96 2.45
ARN -11.61 0.84 1.51 0.94 2.43 0.96 2.92
DeepASC -15.94 0.92 2.21 0.97 3.35 0.98 3.76
(a) Engine
(b) Babble
(c) Factory
(d) Speech
Figure 3: Comparison of NMSE (dB) over time for different noise types.
6.2 Speech Cancellation
Table 7 presents the average NMSE values for different ANC algorithms across three speech datasets: TIMIT, LibriSpeech, and WSJ, with speech segments affected by varying levels of nonlinear distortions. As observed in the noise cancellation case, in speech cancellation, the non-deep learning methods—FxLMS and THF-FxLMS—demonstrate suboptimal performance compared to deep learning-based approaches. Among the deep learning methods, DeepASC achieves the best overall results, surpassing the other algorithms significantly.
In the case without nonlinear distortions ($\eta^{2}=\infty$), DeepASC shows improvements over ARN by 6.13 dB, 4.78 dB, and 5.95 dB for the TIMIT, LibriSpeech, and WSJ datasets, respectively. In the presence of moderate nonlinear distortions ($\eta^{2}=0.5$), DeepASC continues to outperform ARN, with improvements of 6.18 dB for TIMIT, 4.34 dB for LibriSpeech, and 5.99 dB for WSJ. Under more severe nonlinear distortions ($\eta^{2}=0.1$), DeepASC maintains its superior performance, with enhancements of $5.97$dB, $2.46$dB, and $5.81$dB for TIMIT, LibriSpeech, and WSJ datasets, respectively. Figure 4 compares the power spectra and spectrograms of different ANC methods applied to a speech signal. DeepASC achieves significantly better noise suppression across all frequencies, particularly in the high-frequency range - DeepASC outperforming DeepANC and ARN. As shown in Figure 3(d), DeepASC consistently yields lower NMSE across nearly all time steps,
However, it is worth noting that the standard NMSE metric may not fully reflect DeepASC’s effectiveness in speech-active scenarios, due to its sensitivity to silent intervals within speech recordings. Examination of the spectrograms in Figure 4 and perceptual listening tests of the error signals highlight a qualitative performance gap between DeepASC and baseline methods that is not entirely captured by NMSE alone. To address this, we performed an evaluation using VAD-based masking to isolate speech-active regions, computing $\text{NMSE}[\text{VAD}(\textbf{P}\ast\textbf{x}),\text{VAD}(\textbf{S}\ast\{f_{LS}\{\textbf{y}\})]$ (see Appendix D for VAD mask examples and details). The results in Table 2 for the $\eta^{2}=0.5$ case reveal a more pronounced advantage: DeepASC outperforms the alternative methods by 8.1,dB on TIMIT, 7.63,dB on WSJ, and 5.03,dB on LibriSpeech.
@@ 6.2 Speech Cancellation
Table 7: Average NMSE ($\downarrow$) in dB for DeepASC and other algorithms across various speech datasets and nonlinear distortions. Lower values indicate better performance.
@@
Figure 4: Spectrograms and Power Spectra of Speech Signal (00da010c from WSJ) using Different ANC methods without nonlinear distortions ($\eta^{2}=\infty$)
Table 8: Average NMSE ($\downarrow$) in dB for DeepASC and other algorithms under real-time constraints.
Method Runtime (S) Engine ($\downarrow$) Babble ($\downarrow$) TIMIT ($\downarrow$) LibriSpeech ($\downarrow$) WSJ ($\downarrow$)
ARN 0.0116 -10.73 -9.37 -12.40 -6.32 -6.61
DeepANC 0.0125 -11.13 -9.84 -12.18 -7.61 -8.33
DeepASC 0.0136 -16.56 -13.14 -14.81 -12.52 -13.40
6.3 Real-World Simulation
We expanded our investigation to assess the performance of our method in real-world settings, testing it across various simulation scenarios. This was necessary because the fixed task acoustic setup, which relies on the image method, has limitations regarding generalizability and real-world performance. We utilized the dataset from [76], which includes acoustic paths from 23 individuals, measured in the real world and encompassing both primary and secondary paths. We applied DeepASC, along with baseline approaches, to the updated simulation conditions, evaluating their performance using Factory and Babble noise from the NoiseX-92 dataset and speech samples from the WSJ dataset. The results in Table 2 present the average NMSE across these categories. The results demonstrate that DeepASC consistently outperforms the alternative methods, achieving improvements of $2.80$dB in the Factory noise, $2.70$dB in the Babble noise, and $1.53$dB on the WSJ dataset.
6.4 Runtime Analysis
Real-time performance is critical in ANC systems. To ensure compliance with the causality constraint, we adopt a future-frame prediction strategy as employed in [6, 41]. Let $T_{p}$ and $T_{s}$ denote the acoustic delays of the primary and secondary paths, respectively, and $T_{\text{ANC}}$ be the algorithm’s processing time. The system must satisfy the constraint $T_{\text{ANC}}<T_{p}-T_{s}$, which evaluates in our setting to $T_{\text{ANC}}<\frac{2}{343}-\frac{0.5}{343}=0.0043$s.
To meet this, DeepASC is optimized for edge deployment using a single S-band, NOAS optimization, and future prediction. Experimental results summarized in Table 8 (all measured using Nvidia H100 80GB GPU) show that while DeepANC and ARN yield slightly lower clock-time latencies (up to 2.6 ms), the difference is negligible due to all models relying on future frame prediction within a 0.01s window (160 samples at 16 kHz). It is demonstrated that DeepASC meets real-time constraints while achieving superior ANC performance related to other methods by NMSE margins of 5.43, 3.3, 2.41, 5.07, and 4.91 dB on engine noise, babble noise, TIMIT, WSJ, and LibriSpeech, respectively.
@@ 6.4 Runtime Analysis
The computational complexity of the models was additionally assessed by comparing their FLOPs, averaged across 20 three-second samples from the Noisex-92 dataset, as presented in Table LABEL:tab:flops_nmse. The single-band, small variant of DeepASC demonstrated exceptional efficiency, requiring only 2.862G FLOPs while consistently surpassing the performance of the other models. This highlights its superior balance between computational cost and effectiveness.
Table 9: Performance comparison across datasets for different DeepASC variants.
Method Factory ($\downarrow$) TIMIT ($\downarrow$) LibriSpeech ($\downarrow$) WSJ ($\downarrow$)
DeepASC (31.9M) -15.94 -16.36 -16.95 -15.32
DeepASC-1L Band (32M) -15.72 -15.95 -16.64 -15.32
DeepASC-LSTM (33.4M) -11.53 -13.33 -13.92 -13.28
DeepASC-Transformer (38M) -14.93 -14.01 -15.2 -13.76
DeepASC-No-Dual (31.9M) -11.59 -12.36 -13.07 -12.3
DeepASC-No-Masking (31.9M) -3.36 7.43 -6.82 1.37
6.5 Ablation Study
We conducted ablation experiments to evaluate the impact of key components in DeepASC, including the Masking mechanism, Mamba layer, multi-band processing and dual-path structure. For the Mamba layer, we replaced it with either a Transformer or LSTM in the masknet. The Transformer-based model used 12 layers for the full-band and 6 layers for small bands, with 2 blocks per layer (d_model=256, 4 heads, d_ffn=1024). The LSTM variant employed the same depth and block structure, each block comprising two LSTM layers (hidden_dim=256). The original DeepASC uses Mamba blocks (d_model=256, ssm_dim=16, mamba_conv=4) with 16 layers for the full-band and 8 for small bands. For the masking mechanism, we eliminated it entirely, allowing direct anti-signal prediction. To test multi-band processing, we trained a single-band model with an equivalent parameter count to the 3-band model. To isolate the dual-path structure, we removed it while preserving activation shapes. Unless noted otherwise, all models used the 3-band configuration (one full-band M and two sub-band S paths), without NOAS optimization.
Results summarized in Table 9 (with $\gamma^{2}=0.5$) show the masking mechanism is crucial—its removal degrades performance by at least 10.13 dB (LibriSpeech). The Mamba block significantly outperforms alternatives: while the Transformer performs comparably, it lags by at least 1.01 dB (Factory); the LSTM model performs worse across all datasets. Finally, although the single-band model matches the 3-band setup on WSJ, the 3-band variant consistently outperforms it elsewhere, confirming the effectiveness of multi-band processing beyond parameter scaling. These findings support our architectural decisions and demonstrate the efficacy of DeepASC. Appendix B provides a further ablation study, focusing on the importance of NOAS optimization.
7 Conclusion and Limitations
This paper introduced a novel ASC method based on the Mamba-Masking architecture. By decomposing and transforming the encoded signal through the masking, our model enhances anti-signal generation and phase alignment, leading to more effective cancellation. Combined with an optimization-based loss (NOAS), the approach achieves near-optimal performance, improving ANC and ASC by 7.2 dB and 6.2 dB respectively over state-of-the-art baselines on voice signals. These results underscore the Mamba-Masking Network’s capacity to manage diverse frequencies and real-world acoustic conditions, where conventional models often under-perform. Despite empirical gains from components like Mamba layers and NOAS, a rigorous theoretical justification for their effectiveness remains an open question. Additionally, we have yet to fully exploit the Mamba architecture’s long-context modeling capabilities. Overall, our framework addresses key limitations in current ANC systems, and opens new directions for advanced audio cancellation technologies.
Appendix
Appendix A NOAS Design Choices & Motivation
As previously discussed, the optimization process is conducted in the S-projected space rather than directly within the domain of the canceling signal itself—that is, by minimizing $\text{NMSE}\left[\mathbf{S}\ast\mathbf{y}^{*},\mathbf{S}\ast\mathbf{y}\right]$ instead of $\text{NMSE}\left[\mathbf{y}^{*},\mathbf{y}\right]$. This projection via S utilizes prior knowledge captured during the initial training phase, specifically the temporal dependencies embedded in the structure of $\mathbf{y}$. For the sake of clarity, we will use the mean squared error (MSE) as the distance metric. To illustrate this, consider the example depicted in Figure 5. In this illustrative case, we assume both P and S are defined as simple averaging filters (e.g., $[0.5,0.5]$ for a two-dimensional signal). Let $\mathbf{y}^{*}$ denote the optimal anti-noise signal such that $\mathbf{P}\ast\mathbf{x}=\mathbf{S}\ast\mathbf{y}^{*}=0$. Additionally, for the model’s output signal $\mathbf{y}$, we have $\mathbf{P}\ast\mathbf{x}=\mathbf{S}\ast\mathbf{y}=0$, which indicates that $\mathbf{y}$ is already optimal in the projected space. However, if we were to directly optimize in the native domain of $\mathbf{y}$ without regard to the projection, the resulting estimate $\mathbf{y}^{\prime}$—although potentially closer to $\mathbf{y}^{*}$—might lead to suboptimal performance since $\mathbf{P}\ast\mathbf{x}\neq\mathbf{S}\ast\mathbf{y}^{\prime}$.
Figure 5: S-projection importance visualization for NOAS optimization.
This behavior is attributed to the properties of convolution with a fixed filter (in this case, S), which does not constitute an isometry and thus fails to preserve distances in the original space. As such, optimization in the S-projected space more faithfully reflects the desired performance criterion.
This conceptual rationale, together with the previously stated motivation, is further substantiated by the provided measurements in Table 1. In particular, the NMSE values were consistently lowest between $\left[\mathbf{S}\ast\mathbf{y}^{*},\mathbf{S}\ast\mathbf{y}\right]$, lending strong support to the claim that $\mathbf{S}\ast\mathbf{y}^{*}$ constitutes a feasible optimization target for $\mathbf{S}\ast\mathbf{y}$. Additionally, a noteworthy observation arises following the NOAS optimization: the NMSE between $\left[\mathbf{P}\ast\mathbf{x},\mathbf{S}\ast\mathbf{y}\right]$ is significantly reduced by 1.07 dB. This empirical finding challenges the notion that NOAS merely functions as a regularization term, instead indicating that it plays a more active role in enhancing the quality of the learned representations.
Appendix B Ablation Study - 2nd Part
@@
To assess the contributions of the NOAS fine-tuning optimization in our method, we conducted an ablation study focusing on multiband processing, band size (small vs. medium), and the impact of NOAS optimization on them. Table 10 presents the results of this analysis, reporting the NMSE across four datasets: Factory, TIMIT, LibriSpeech, and WSJ, all evaluated under nonlinear distortion conditions ($\eta^{2}$ = 0.5).
@@
In our notation, "+ S - Multiband - NOAS" refers to a small band configuration (8 mamba layers) without multiband processing or NOAS optimization, while "+ S - Multiband + NOAS" refers to the same small band architecture with NOAS optimization applied. Similarly, "+ M - Multiband - NOAS" represents a medium band configuration (16 mamba layers) without NOAS, and "+ M - Multiband + NOAS" applies NOAS optimization to the same medium band model. The "+ MultiBand - NOAS" is defined as a configuration that employs one full medium band and two small sub-bands without NOAS optimization applied, whereas the Full Method is defined as the same configuration with NOAS optimization applied.
@@
All models were initially trained using the ANC loss function defined in Eq. 8. Configurations with "+ NOAS" were fine-tuned using NOAS optimization, whereas configurations with "- NOAS" were trained exclusively using the ANC loss in Eq. 8. The results demonstrate that the removal of NOAS optimization consistently degrades performance across all datasets. For instance, on the Factory dataset, applying NOAS optimization to the small band model leads to a performance improvement of $0.73$dB, while the medium band model shows a larger improvement of $0.90$dB. This trend holds across the other datasets, reinforcing the crucial role of NOAS optimization in enhancing model performance. Multiband processing further improves the overall effectiveness of DeepASC. The Full Method consistently outperforms the "+ Multiband - NOAS" configuration, with gains of $0.29$dB, $0.09$dB, $0.13$dB, and $0.15$dB on the Factory, TIMIT, LibriSpeech, and WSJ datasets, respectively. Interestingly, the performance of the "+ M - Multiband + NOAS" configuration is higher than that of the "+ Multiband - NOAS" variant by $0.15$dB. This indicates that while multiband processing is valuable, the choice of band size plays a significant role in the model’s performance, with larger band sizes, particularly when combined with NOAS, yielding the best results.
Appendix C Model Analysis
The number of frequency bands in the DeepASC architecture is a critical hyperparameter affecting performance. Table 11 compares DeepASC’s performance across different band configurations for the Factory noise, TIMIT, LibriSpeech, and WSJ datasets, with $\eta^{2}=0.5$. The "1-band" models use a single full band, while the "3-band" and "4-band" models incorporate one medium band with two and three smaller sub-bands, respectively. A 2-band model, which would require two full bands, was excluded as it falls outside the intended design of DeepASC.
Table 11: Average NMSE ($\downarrow$) in dB of our method (w/o NOAS) for Noise and Speech using different number of bands, with nonlinear distortion of $\eta^{2}=0.5$.
DeepASC (small) 1 -13.46 -14.26 -14.88 -13.22
DeepASC (medium) 1 -15.19 -15.82 -16.56 -14.86
DeepASC 3 -15.94 -16.36 -16.95 -15.32
DeepASC 4 -16.52 -16.55 -17.41 -15.84
Table 12: Comparison of different deep learning based ANC methods based on parameter size.
Models #Params NMSE ($\downarrow$)
Deep-ANC 8.8M -10.69
ARN 15.9M -11.61
DeepASC, 1 Band, S 8.0M -13.46
DeepASC, 1 Band, M 15.8M -15.19
DeepASC, 3 Bands 31.9M -15.94
DeepASC, 4 Bands 40.0M -16.52
Model Size Comparison. Table 12 compares model size and performance, with NMSE evaluated on factory noise under nonlinear distortion of $\eta=0.5$. DeepASC variants in this comparison are without NOAS optimization. The results indicate that even the smallest DeepASC configuration (1-band, small) outperforms the ARN architecture by $1.85$ dB, despite using only half the parameters ($8.0$M vs. $15.9$M). This is a significant outcome given the critical importance of model size in real-time ANC applications where latency is critical.
Appendix D VAD Masks Visualization
We provide additional visualizations and a detailed explanation of the VAD mask employed in our proposed method. Specifically, the audio signals were segmented using a window length of 256 samples with an overlap of 128 samples. The energy threshold for the VAD was set to 10% of the maximum energy observed within the corresponding speech signal. Frames with energy values below this threshold were marked as inactive (i.e., masked). Figure 6 presents the VAD masks applied to 9 distinct speech samples. In each subplot, the red line indicates the binary VAD mask. Segments where the mask is zero correspond to suppressed (i.e., nulled) portions of the signal, whereas segments with a non-zero mask retain the original signal content unaltered.
(a) Sample 020o030d
(b) Sample 020o031e
(c) Sample 020o0318
(d) Sample 203c0b0k
(e) Sample 203c0b12
(f) Sample 203o0b16
(g) Sample 400e080j
(h) Sample 400r0712
(i) Sample 400r0705
Figure 6: Visualization of VAD masks applied to nine different speech signals from WSJ dataset. Each subplot shows the energy contour with the overlaid red VAD mask.